# Time-To-Inconsistency: A Survival Analysis of Large Language Model Robustness to Adversarial Attacks

### v1

@@ 4.1 Data
We conduct a comprehensive robustness evaluation using the MT-Consistency benchmark Li et al. (2025a), which provides a systematic framework for assessing LLM consistency across multi-turn adversarial interactions. Our analysis covers 9 state-of-the-art LLMs across 8 turns of adversarial interactions, analyzing over 36,000 individual model responses.
Benchmark Protocol: Following the MT-Consistency framework, we employ a strict consistency criterion where only conversations with correct initial responses (round 0) are included. Failure is defined as any deviation from the initial correct response in subsequent adversarial rounds, providing a stringent test of model robustness under sustained pressure.
Dataset Composition: The benchmark contains 700 carefully selected questions spanning 39 individual academic subjects across multiple difficulty levels (Elementary, High School, College, Professional). To enable systematic analysis of domain-specific vulnerabilities, we implement a theoretically-motivated subject clustering approach that groups the 39 individual subjects into 7 coherent thematic domains: STEM (11 subjects), Medical Health (8 subjects), Social Sciences (4 subjects), Humanities (6 subjects), Business Economics (5 subjects), Law Legal (3 subjects), and General Knowledge (2 subjects). Complete subject-to-cluster mappings are provided in Appendix A.
This clustering enables both fine-grained subject-level analysis and broader domain-level robustness assessment, revealing patterns that would be obscured in either purely individual-subject or overly-aggregated analyses.
Adversarial Interaction Design: Each conversation consists of an initial question followed by up to 8 systematically designed adversarial follow-up prompts. These prompts are specifically crafted to induce semantic drift and test model consistency, representing 8 distinct adversarial attack patterns that challenge different aspects of model reasoning and memory: Closed-ended (C), Open-ended (O), Misleading (M), Emotional Appeal (EmA), Impolite Tone (IT), Expert Appeal (ExA), Consensus Appeal (CA), and False Agreement (FA). See complete prompt templates in Appendix B.
These adversarial strategies target different psychological and cognitive vulnerabilities, from simple uncertainty induction (C) to sophisticated social pressure tactics (CA, ExA) and deceptive agreement patterns (FA). The diversity of attack vectors ensures comprehensive evaluation of model robustness across multiple dimensions of adversarial pressure.
@@ 4.2 Evaluation Metrics
We assess model performance using a combination of metrics tailored to different aspects of survival prediction.
Discrimination and Calibration: We evaluate all models on two primary metrics: Harrell’s concordance index (C-index) for discrimination and the Integrated Brier Score (IBS) for overall predictive accuracy. The C-index measures a model’s ability to correctly rank the survival times of pairs of conversations, while the IBS measures the mean squared error between predicted survival probabilities and observed outcomes over time, providing a comprehensive assessment of both discrimination and calibration.
Model Selection and Tuning: For model selection among the Cox and AFT specifications, we use the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC). This helps compare, for instance, the baseline versus the interaction Cox model, or the relative fit of different distributional assumptions for the AFT models. For the Random Survival Forest, hyperparameters are tuned by optimizing the out-of-bag (OOB) C-index and IBS.
@@ 4.3 Experiment Setup
To ensure an unbiased assessment of generalization, we partition the dataset into a training set (80%) and a held-out test set (20%). The split is performed at the conversation level, ensuring that all turns from a single dialogue reside in the same set to prevent data leakage. All model training, hyperparameter tuning, and feature selection are conducted exclusively on the training data. The final predictive performance of the models is then evaluated on the untouched 20% test set.
@@ 5.1 Overall Model Performance
Table 1: Comprehensive Model Performance on Test Set
Model Paradigm Features C-index IBS
@@
5.2 Calibration Analysis
@@
Table 2: Brier Score Analysis by Conversation Round
@@
The calibration analysis reveals that AFT models demonstrate remarkable improvement in later conversation rounds, with Brier scores approaching zero by round 8. This pattern suggests that parametric models capture the accelerating nature of conversation degradation more effectively than proportional hazards models.
5.3 Proportional Hazards Assumption Validation
Our Schoenfeld residuals analysis for Cox models reveals systematic violations of the proportional hazards assumption, particularly for semantic drift features. Table 3 summarizes the statistical tests.
Table 3: Proportional Hazards Assumption Test Results (Schoenfeld Residuals)
Prompt-to-Prompt Drift 0.032 0.021 Yes Time-varying effect
Context-to-Prompt Drift 0.067 0.045 Marginal Slight violation
Cumulative Drift 0.156 0.089 No Assumption holds
Model Interactions 0.003 Yes Strong violation
Length Features 0.234 0.187 No Assumption holds
Repetition Metrics 0.421 0.356 No Assumption holds
The systematic violations of proportional hazards assumptions for key semantic drift features ($p<0.05$) provide strong empirical justification for our multi-paradigm modeling approach. These violations suggest that the effect of semantic drift on failure hazard changes over time, supporting the use of AFT models that naturally accommodate time-varying effects.
5.4 Feature Importance and Risk Factor Analysis
Figure 1 summarizes model–drift interactions from the advanced Cox PH (HRs; dashed line == neutral effect). Three patterns emerge:
Figure 1: Semantic Drift Effects on Failure Risk: p2p drift consistently increases failure risk (HR >> 2), while cumulative drift often provides protection (HR << 1), suggesting sophisticated adaptation mechanisms across models.
(1) Prompt-to-prompt (p2p) drift is catastrophic. Acute turn-to-turn shifts (p2p) are the dominant failure driver across all LLMs, with large hazard ratios: $\mathrm{HR}\!\approx\!1.2$ (Gemini), $2.3$ (gpt_oss), $4.6$ (gpt_4o), $3.2$ (llama_4), $2.8$ (qwen). Thus, even small immediate semantic jumps sharply elevate next-turn failure risk, especially for gpt_4o and llama_4.
(2) Context-to-prompt drift poses moderate risk. Deviations from the running context (c2p) are harmful but weaker than p2p, with $\mathrm{HR}\!\approx\!1.1$$3.1$ across models. This suggests failures are more sensitive to sudden shocks than to gradual divergence from the cumulative context.
(3) Cumulative drift is protective. Counterintuitively, higher accumulated drift (cum) is associated with lower risk (all $\mathrm{HR}\!<\!1$, roughly 0.20.20.80.8). A plausible interpretation is adaptation: once a conversation has survived several shifts, the model may stabilize to the evolving topic (or adversarial pressure decays), reducing incremental hazard.
Overall, p2p emerges as the principal actionable risk factor (acute shocks), c2p as a secondary risk (context divergence), and cumulative drift as a resilience marker. These effects are consistent in sign across models but vary in magnitude, revealing distinct vulnerability profiles.
@@ 5.5 Temporal Failure Patterns
All models demonstrate statistically significant risk stratification ($p<0.001$), with hazard ratios ranging from 1.87 (RSF) to 2.67 (Cox Advanced). The consistent pattern across modeling paradigms provides robust evidence for the predictive validity of our semantic drift features.
@@ 6 Discussion
Our findings offer a new perspective on the robustness of Large Language Models in multi-turn dialogues, shifting the focus from static, single-turn accuracy to the temporal dynamics of conversational failure. This work demonstrates that the path to inconsistency is not random but a predictable process driven by the nature of the semantic drift. The central discovery is the starkly different roles of abrupt versus gradual drift. We found that abrupt, P2P semantic shifts act as catastrophic shocks that dramatically increase the immediate risk of failure. Conversely, gradual, cumulative drift over a conversation is paradoxically protective, suggesting that models can adapt to and even become more robust within a coherently evolving dialogue. This challenges the conventional wisdom that all deviation from an initial topic is detrimental, indicating instead that the velocity of semantic change is a more critical determinant of conversational integrity than the total distance traveled.
The superior performance of Accelerated Failure Time (AFT) models is not merely a statistical artifact but a direct reflection of the underlying failure process. Our analysis confirmed that the proportional hazards assumption—the foundation of simpler Cox models—is systematically violated for key drift features. This means the risk of failure is not constant; it accelerates as a conversation progresses under adversarial pressure. AFT models excel precisely because they are built to capture this time-varying nature of risk, explaining their superior calibration and predictive accuracy, especially in the crucial later rounds of a dialogue. This methodological insight is critical: to accurately predict and understand LLM failure, we must employ analytical tools that respect the dynamic, non-constant nature of the hazard.
These insights have immediate practical applications for the entire LLM lifecycle. The dominance of p2p drift as a failure catalyst provides a clear mandate for developing real-time monitoring and early warning systems tuned to detect these acute conversational shocks. By using a lightweight AFT model, a system can move beyond post-hoc analysis to proactive intervention. Such a monitor could identify at-risk conversations with high discriminative accuracy (C-index up to 0.874) and provide exceptionally well-calibrated failure probabilities (IBS <0.18<0.18). This enables sophisticated risk stratification in production, allowing for dynamic resource allocation, graceful topic changes, or timely hand-offs to human agents before a user’s trust is irrevocably broken.
@@ 7 Conclusion
By reframing multi-turn conversational failure as a time-to-event process, this work establishes a powerful new paradigm for evaluating LLM robustness. We demonstrated that the path to inconsistency is a predictable process governed by the velocity of semantic drift, where abrupt conversational shocks are catastrophic and gradual topical evolution is a marker of resilience. Methodologically, we provided conclusive evidence that the risk of LLM failure is non-constant, a critical finding that validates the superior performance of Accelerated Failure Time models and highlights the limitations of traditional proportional hazards assumptions in this domain. Ultimately, our survival analysis framework provides the tools to move beyond static, post-hoc benchmarks and toward the dynamic, real-time monitoring of conversational health, paving the way for the development of more resilient and reliable AI agents.
Appendix A Subject Domain Clustering Details
A.1 Complete Subject-to-Cluster Mappings

### v2

@@ 4.1 Data
We conduct our study on the MT-Consistency li-etal-2025-firm, which systematically probes LLM consistency under adversarial multi-turn interactions. Each conversation is built from a base question followed by up to 8 adversarial follow-ups; we adopt this 8-turn horizon in all experiments.
Questions and subjects.
The benchmark contains 700 questions spanning 39 academic subjects and four difficulty bands (Elementary, High School, College, Professional). To support both fine-grained and domain-level analysis, we group the 39 subjects into 7 thematic clusters: STEM (11 subjects), Medical Health (8), Social Sciences (4), Humanities (6), Business Economics (5), Law/Legal (3), and General Knowledge (2). The complete mapping is provided in Appendix B.
Models.
We evaluate nine state-of-the-art LLMs: Claude 3.5 Sonnet, DeepSeek R1, GPT-4o, an open-weight 120B GPT-style model (gpt_oss_120B), Llama 3.3 70B, Llama 4 Maverick, Gemini 2.5, Mistral Large, and Qwen 3. For each base question, all nine models are evaluated under the same adversarial prompt templates, yielding a matched set of multi-turn trajectories. Unless otherwise stated, we pool conversations from all models into a single dataset and include model identity $M_{i}$ as a covariate in $\mathbf{X}_{i,t}$. After filtering for initially correct answers, the resulting corpus comprises 36,951 turns across all models.
Adversarial interaction design.
Each conversation consists of an initial question followed by up to 8 systematically designed adversarial follow-up prompts. These prompts are crafted to induce semantic drift and test consistency, covering 8 attack patterns: Closed-ended (C), Open-ended (O), Misleading (M), Emotional Appeal (EmA), Impolite Tone (IT), Expert Appeal (ExA), Consensus Appeal (CA), and False Agreement (FA). Full templates are given in Appendix C. Together, these strategies range from mild uncertainty induction to strong social-pressure tactics, providing a diverse stress test for multi-turn robustness.
@@ 4.2 Evaluation Metrics
We evaluate survival models along two complementary dimensions:
Discrimination.
We use Harrell’s concordance index (C-index) to measure how well a model ranks conversations by time-to-inconsistency. A C-index of 0.5 corresponds to random ordering; higher values indicate better ability to assign higher risk to conversations that fail earlier.
Calibration and overall accuracy.
We compute Brier scores at each turn $t=1,\dots,8$ and report the Integrated Brier Score (IBS), which averages the Brier score over time. The IBS captures both discrimination and calibration of predicted survival probabilities $\hat{S}_{i}(t)$, with lower values indicating more accurate and better-calibrated risk predictions.
@@ 4.3 Experiment Setup
We split conversations at the conversation level into an 80% training pool and a 20% held-out test set, stratified by model and subject cluster to preserve their marginal distributions. All test-set metrics (C-index, Brier scores, IBS) are computed once on this 20% and are not used for model selection or hyperparameter tuning.
Within the 80% training pool, we perform 5-fold cross-validation over conversations to tune hyperparameters and select model variants:
  • Cox models: we treat the strength of $\ell_{2}$ regularization on drift–model interaction terms and the choice between a baseline-only and an interaction specification as hyperparameters. We select these using 5-fold cross-validated IBS on the training pool, with C-index as a secondary tie-breaking criterion.
  • AFT models: we consider Weibull, log-normal, and log-logistic baseline distributions, and jointly tune the distribution family and $\ell_{2}$ regularization strength. The selected configuration is the one that achieves the best 5-fold cross-validated IBS on the training pool.
  • RSF: we tune the number of trees, maximum depth, and the number of variables tried at each split (mtry), again using 5-fold cross-validated IBS.
This procedure ensures that all hyperparameters and model choices are determined using only the training pool (via internal cross-validation), and the test set is used exactly once for final evaluation. The full search grids and the selected configurations for each model are reported in Appendix E.
@@ 5.1 Overall Model Performance
Table 1: Model performance on the held-out test set. Higher C-index and lower IBS are better.
Model Paradigm # of covariates C-index IBS
@@
5.2 Calibration analysis over Turns
@@
Table 2: Brier score by conversation round on the test set. Lower is better.
@@
Cox models’ Brier scores increase monotonically and remain relatively high in later rounds, reflecting overconfident survival estimates as adversarial pressure accumulates. In contrast, AFT models’ Brier scores flatten and then decrease toward the end of the horizon (rounds 7–8), indicating that they better capture the accelerating nature of failure risk in this adversarial setting. RSF tracks the AFT models reasonably well but with slightly higher Brier scores at later turns.
Taken together with the C-index results, this suggests that parametric AFT assumptions provide a good approximation to the true time-to-inconsistency process in MT-Consistency, especially when modeling the shape of risk over turns.
Proportional hazards check.
We also verify the proportional hazards (PH) assumption for the Cox models using Schoenfeld residual tests. Key semantic drift covariates, especially prompt-to-prompt drift, show clear departures from PH, while length and most subject/difficulty indicators do not. Full p-values and diagnostics are reported in Appendix D.
5.3 Robustness of Feature Importance Analysis
To ensure our insights are not artifacts of model misspecification, we cross-verified Cox PH results against the AFT model, which does not rely on the PH assumption. Figure 1 presents the comparison. Note the inverse relationship required for consistency: a high Hazard Ratio ($\mathrm{HR}>1$) in Cox corresponds to a low Acceleration Factor ($\mathrm{AF}<1$, implying shortened survival time) in AFT.
Figure 1: Robustness Check: Cox Hazard Ratios vs. AFT Acceleration Factors. The models show strong directional agreement. P2P drift (Red) consistently increases risk ($\mathrm{HR}>1,\mathrm{AF}<1$), while Cumulative drift (Green) is consistently protective ($\mathrm{HR}<1,\mathrm{AF}>1$).
(1) Prompt-to-Prompt (p2p) drift is undeniably catastrophic. Despite the PH violation, both models identify acute semantic shifts as the dominant failure driver. The Cox model estimates severe risk (e.g., GPT-4o $\mathrm{HR}\approx 4.7$), which is corroborated by the AFT model estimating a drastic reduction in expected conversation length (GPT-4o $\mathrm{AF}\approx 0.15$). This confirms that immediate semantic jumps destabilize the model regardless of the temporal distribution assumptions.
(2) Cumulative drift is genuinely protective. One of our main insights—that accumulated drift is protective—holds true under the AFT framework. While Cox shows reduced hazard ($\mathrm{HR}<1$), the AFT model estimates a time expansion factor of $1.4\times$ to $2.6\times$ across models. This validation suggests that the protective effect is not a statistical artifact: as conversations progress and ”survive” early turns, models effectively adapt to the drifting context.
(3) Consistency across model architectures. The concordance between Cox and AFT results validates the stability of our feature importance hierarchy: $\text{P2P}>\text{C2P}>\text{Cumulative}$ (Protective). Crucially, these qualitative patterns persist across both model specifications, demonstrating that our primary insights are robust to the Proportional Hazards assumption violation.
@@ 5.4 Temporal Failure Patterns
Across all modeling paradigms, high-risk conversations terminate much earlier than low-risk ones: median survival times drop from roughly 8 turns (censored at the horizon) in the low-risk group to about 4–4.5 turns in the high-risk group. Log-rank tests strongly reject equality of survival curves ($p<0.001$ in all cases), and hazard ratios between high- and low-risk strata range from 1.87 (RSF) to 2.67 (Cox advanced). This confirms that the features used by our models—particularly the drift covariates—support meaningful risk stratification: they are not only predictive at a single horizon, but also separate conversations into trajectories with qualitatively different robustness under sustained adversarial pressure.
5.5 Retrospective Risk Monitoring with AFT
Finally, we investigate the operational utility of our best-performing AFT model as a real-time safeguard. While predictive accuracy (C-index) is important, a practical monitor must offer actionable lead time while minimizing alert fatigue. Rather than using a static failure time prediction, we compute a Conditional Failure Probability (CFP) over a rolling horizon $\tau$. At any turn $t$, given that the conversation is currently consistent ($T>t$), the probability of failure occurring within the next $\tau{=}2$ turns is $\text{Risk}_{i}(t,\tau)=1-\frac{\hat{S}_{i}(t+\tau)}{\hat{S}_{i}(t)}.$ This metric dynamically updates based on the accumulated hazard, and we trigger an alert when this risk exceeds a threshold $\lambda$ optimized for F1 during training.
Table 4: Behavior of the AFT-based risk monitor and a drift-threshold baseline on the test set. “% alerted” is the fraction of conversations in which at least one alert is raised before failure or censoring. “Alerts / conv.” is the mean number of alerts per conversation within each group. “First-alert round” and “Failure round” are means over conversations in the corresponding group (“–” where no failure occurs).
Group Method % alerted Alerts / conv. First-alert round Failure round
All (140) AFT (ours) 55% 1.1 4.0
Drift baseline 51% 1.3 4.0
Failing (88) AFT (ours) 76% 1.4 3.3 5.7
Drift baseline 62% 1.6 3.9 5.7
Censored (52) AFT (ours) 19% 0.5 5.2
Drift baseline 32% 1.2 4.2
Applying this AFT-based monitor to the held-out test set demonstrates highly effective intervention capabilities (Table 4). The monitor successfully triggers an alert for 76% of failing conversations before the inconsistency occurs, and among these correctly warned dialogues the system provides a median lead time of 2 turns (mean 2.3 turns) between the first warning and the actual event. This indicates that the model detects precursors of failure—specifically the accelerating hazard induced by semantic drift—well within a viable intervention window. At the same time, the system remains operationally selective: only 19% of censored (safe) conversations ever trigger an alert within the 8-turn window. Alert density also differs sharply by outcome: the monitor raises on average 1.4 alerts per failing dialogue but only 0.5 per censored dialogue (about 1.1 per conversation overall). The drift-threshold baseline, in contrast, alerts fewer failing conversations (62%) while generating more noise on safe ones (32% censored alerted, 1.2 alerts per censored dialogue). It also tends to fire later in failing dialogues (mean first-alert round 3.9 vs. 3.3 for AFT).
@@ 6 Discussion
Our findings offer a new perspective on the robustness of Large Language Models in multi-turn dialogues, shifting the focus from static, single-turn accuracy to the temporal dynamics of conversational failure. This work demonstrates that the path to inconsistency is not random but a predictable process driven by the nature of the semantic drift. The central discovery is the starkly different roles of abrupt versus gradual drift. We found that abrupt, prompt-to-prompt (P2P) shifts act as catastrophic shocks that dramatically increase the immediate risk of failure. Conversely, gradual, cumulative drift over a conversation is paradoxically protective, suggesting that models can adapt to and even become more robust within a coherently evolving dialogue. This challenges the conventional wisdom that all deviation from an initial topic is detrimental, indicating instead that the velocity of semantic change is a more critical determinant of conversational integrity than the total distance traveled.
Methodologically, our results highlight the importance of choosing survival models whose assumptions match the underlying failure process. The proportional hazards (PH) checks in Appendix D indicate that key semantic drift covariates, especially P2P drift, violate the PH assumption: their effects on hazard are not constant over turns. This aligns with the intuition that adversarial pressure reshapes risk as conversations progress. In this setting, Cox models remain useful as descriptive tools—e.g., for summarizing average hazard ratios—but are mis-specified as fully generative models of time-to-inconsistency. In contrast, parametric Accelerated Failure Time (AFT) models explicitly act on the time scale and are better aligned with an accelerating risk profile. This helps explain their superior calibration and predictive accuracy, especially in the crucial later rounds of a dialogue. This methodological insight is critical: to accurately predict and understand LLM failure, we must employ analytical tools that respect the dynamic, non-constant nature of the hazard.
Finally, our retrospective monitoring experiment illustrates that survival models are not only analytically insightful but also operationally useful. A lightweight Weibull AFT model, tuned only on training data, attains high discriminative accuracy (test C-index up to 0.874, IBS <0.18<0.18) and can be converted into a simple turn-wise risk score that drives concrete safeguards, with the monitoring results demonstrating that such scores meaningfully anticipate failure rather than merely describing it post hoc. In this sense, survival analysis turns multi-turn robustness from a static summary into an evolving risk signal, opening the door to agents that do not merely fail more slowly, but actively recognize when a dialogue is entering a dangerous regime and adapt their behavior accordingly. This perspective enables more sophisticated risk stratification in deployment, including dynamic allocation of oversight, graceful topic shifts or clarifying questions when risk spikes, and timely hand-offs to human operators before a user’s trust is irrevocably broken.
Limitations
First, all experiments are conducted on MT-Consistency, with one family of adversarial prompt protocols and a maximum horizon of eight turns. While this provides a controlled environment for analysis, it does not cover longer, mixed-initiative dialogues or other adversarial styles (e.g., tool use, or chain-of-thought steering). Second, we treat the first inconsistent answer as a binary event, without distinguishing between qualitatively different failure types (sycophancy, hallucination, instruction misinterpretation, etc.), and we rely on a single embedding model to define semantic drift. Third, our monitoring analysis is purely retrospective: the AFT-based risk scores are evaluated offline and not coupled to real interventions or user outcomes.
These limitations suggest several concrete directions for future work. On the evaluation side, extending time-to-inconsistency analyses to other domains, attack families, and longer horizons would test how general our drift–hazard findings are. On the modeling side, adding richer covariates—such as confidence estimates, response-level features, or error-type labels—could better disentangle failure modes and improve interpretability. On the deployment side, integrating survival-based monitors into real systems with human-in-the-loop interventions and online A/B tests would let us directly measure their impact on safety and trust. Our results provide an initial step, showing that survival analysis can turn static robustness scores into temporally resolved, actionable risk signals.
@@ 7 Conclusion
By reframing multi-turn conversational failure as a time-to-event process, this work establishes a powerful new paradigm for evaluating LLM robustness. We demonstrated that the path to inconsistency is a predictable process governed by the velocity of semantic drift, where abrupt conversational shocks are catastrophic and gradual topical evolution is a marker of resilience. Methodologically, we provided conclusive evidence that the risk of LLM failure is non-constant, a critical finding that validates the superior performance of Accelerated Failure Time models and highlights the limitations of traditional proportional hazards assumptions in this domain. Ultimately, a lightweight Weibull AFT fit can be converted into a simple conditional-failure monitor that issues early warnings for most failing conversations several turns before the first inconsistent answer while keeping false alerts modest. In this way, survival analysis turns multi-turn robustness from a static benchmark into an evolving risk signal, opening the door to conversational agents that not only fail more slowly but also recognize when a dialogue is entering a dangerous regime and adapt or escalate accordingly.
Appendix A AFT Model Families and Closed-Form Survival Functions
For completeness, we summarize the survival and hazard functions for the parametric Accelerated Failure Time (AFT) models used in this work. In all cases, we write
$\log T=\mu+\sigma\,\varepsilon,$
where $\mu=\bm{\theta}^{\top}\mathbf{Z}$ and $\sigma>0$.
Weibull AFT.
If $\varepsilon$ follows an extreme-value distribution, then $T$ has a Weibull distribution with shape $k=1/\sigma$ and scale $\lambda=\exp(\mu)$. The survival and hazard functions are
$S(t)=\exp\!\left\{-\left(\frac{t}{\lambda}\right)^{k}\right\},\qquad h(t)=\frac{k}{\lambda}\left(\frac{t}{\lambda}\right)^{k-1}.$
Log-normal AFT.
If $\varepsilon\sim\mathcal{N}(0,1)$, then $T$ is log-normally distributed with
$S(t)=1-\Phi\!\left(\frac{\ln t-\mu}{\sigma}\right),\qquad h(t)=\frac{f(t)}{S(t)},$
where $f(t)$ is the log-normal density and $\Phi(\cdot)$ is the standard normal CDF.
Log-logistic AFT.
If $\varepsilon$ follows a standard logistic distribution, then $T$ has a log-logistic distribution with shape $k=1/\sigma$ and scale $\lambda=\exp(\mu)$. The survival and hazard functions are
$S(t)=\frac{1}{1+\left(\tfrac{t}{\lambda}\right)^{k}},\qquad h(t)=\frac{(k/\lambda)\left(\tfrac{t}{\lambda}\right)^{k-1}}{1+\left(\tfrac{t}{\lambda}\right)^{k}}.$
In all cases, changes in $\mu$ induced by covariates correspond to multiplicative changes in characteristic times (e.g., medians), which we interpret via acceleration factors in the main text.
Appendix B Subject Domain Clustering Details
B.1 Complete Subject-to-Cluster Mappings